OpenStudy (anonymous):

which choice is the solution to the equation below? (x-3)^2=25 ax=3orx=5 bx=9orx=-2 cx=-2orx=6 dx=1orx=2

OpenStudy (anonymous):

i think its a is it

OpenStudy (anonymous):

A???????????????????????///////

OpenStudy (anonymous):

nope!

OpenStudy (anonymous):

wattgttttttt

OpenStudy (anonymous):

b

OpenStudy (anonymous):

What is B? B. x=9 or x=-2 B. x=8 or x=-2

OpenStudy (anonymous):

STOP GUESSING!

OpenStudy (anonymous):

B. x=9 or x=-2

OpenStudy (anonymous):

What is C? C. x=-2 or x=6 C. x=-2 or x=8

OpenStudy (anonymous):

C. x=-2 or x=6

OpenStudy (anonymous):

THEN NONE OF THE ABOVE!

OpenStudy (anonymous):

wait b is x=8 or x=-2

OpenStudy (anonymous):

Seriously dude. Stop guessing. Work it out. And make sure everything is RIGHT. Stop rushing \[(x-3)^2=25 \]\[x^2-6 x+9 = 25\]\[x^2-6 x-16 = 0\]

OpenStudy (anonymous):

do you KNOW what the answer is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what

OpenStudy (anonymous):

x=8 or x=-2

OpenStudy (anonymous):

YES! FINALLY :)

OpenStudy (anonymous):

thanks xD